OpenStudy (anonymous):

Anyone In consumers math ?? I'm stuck on the test. Help please!! Medal and fan for help

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

helloooooooo @xsavannah_leigh ????

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \ln(consumers~\cdot~math) = \ln(consumers)+ \ln(math)}\)

OpenStudy (solomonzelman):

familiar with natural logarithms ?

OpenStudy (anonymous):

Your parents are buying a house for $187,500. They have a good credit rating, are make ng a 20% down payment, and expect to pay $1,75/month. The interest rate for the mortgage is 4.65%. What must their realized income be before each months and how much interest is paid at the end of rbe second month?

OpenStudy (anonymous):

@RasheedahOwl @ayanna1100